# Explainable Compliance Detection with Multi-Hop Natural Language Inference on Assurance Case Structure

### v1

@@ Abstract
Ensuring complex systems meet regulations typically requires checking the validity of assurance cases through a claim-argument-evidence framework. Some challenges in this process include the complicated nature of legal and technical texts, the need for model explanations, and limited access to assurance case data. We propose a compliance detection approach based on Natural Language Inference (NLI): EXplainable CompLiance detection with Argumentative Inference of Multi-hop reasoning (EXCLAIM). We formulate the claim-argument-evidence structure of an assurance case as a multi-hop inference for explainable and traceable compliance detection. We address the limited number of assurance cases by generating them using large language models (LLMs). We introduce metrics that measure the coverage and structural consistency. We demonstrate the effectiveness of the generated assurance case from GDPR requirements in a multi-hop inference task as a case study. Our results highlight the potential of NLI-based approaches in automating the regulatory compliance process.
@@ 1 Introduction
Previous study (Cejas et al., 2023) uses inherently interpretable rule-based graph matching, but it still struggles with linguistic variability, ambiguous legal phrasing, and scalability to new regulations. A recent study (Azeem and Abualhaija, 2024) explores statistical machine learning and language model methods for the compliance detection task as a text classification problem. However, it does not consider transparency and explainability factors such as model interpretability and reasoning.
To address the transparency and robustness factors of detecting risk and safety, experts use assurance cases (Rhodes et al., 2010; Mansourov and Campara, 2011) as a structured argument approach to audit complex software products, such as those in aircraft or autonomous vehicles, against standard requirements. This is to ensure compliance with legal and technical standards. Assurance case is a systematic collection of arguments and supporting evidence to demonstrate a certain level of confidence that a product meets particular claims related to its requirements (Netkachova et al., 2015; Piovesan and Griffor, 2017). These assurance cases can be created and updated throughout the development lifecycle (Ross et al., 2022).
@@
One possible approach is to use deductive reasoning (Rushby et al., 2015) on the assurance case. Deductive reasoning in the assurance case derives conclusions from structured claims and arguments from evidence (Bloomfield and Bishop, 2009). This reasoning can be framed in the Natural Language Inference (NLI) task (Dagan et al., 2005; Bowman et al., 2015) to determine if a hypothesis follows from a premise. If there is a set of premises, multi-hop inference (Jansen, 2018) is needed for reasoning over interconnected claims and evidence. This makes multi-hop inference suitable for assurance cases that link high-level system properties to low-level evidence, like test results.
This paper proposes EXplainable CompLiance detection with Argumentative Inference of Multi-hop reasoning (EXCLAIM) based on Assurance Case. Figure 2 illustrates our workflow that explains this paper’s contribution, which provides a summary of this paper. From the figure, we use pre-trained transformer (Devlin et al., 2019) models to learn premise-hypothesis pairs for deductive multi-hop reasoning to perform compliance detection.
Specifically, our contributions address challenges in (i) explainable compliance detection and (ii) the lack of assurance case data and advanced reasoning, as follows:
  • We propose compliance detection based on Natural Language Inference. We formulate the claim-argument-evidence as a discourse structure which can be used to develop a multi-hop inference model for explainable tracing of the compliance detection model’s output (Figure 1).
  • We propose evaluation methods to measure the assurance case generated by large language models (LLMs) to validate the reliability of instance-level consistency and structural coherence. We conduct a study by generating Claim Argument Evidence Assurance Case data with open source and proprietary LLMs from GDPR requirements as the context prompt from a previous study (Cejas et al., 2023). We analyse the generated Assurance Case data by using our proposed metrics.
@@ 4.1 GDPR-DPA
We use the compliance detection dataset from the previous study (Azeem and Abualhaija, 2024) for the NLI task on the real-world human-annotated dataset  [^1] https://zenodo.org/records/11047441. We called this dataset as GDPR-DPA from General Data Protection Regulation-Data Processing Agreement (GDPR-DPA). The details of the dataset statistics can be seen in Table 2. The number of instances in training set is larger than original data due to pairing of negative samples between non-matching premise (requirements) and the hypothesis (DPA). We set the sampling rate 0.1 to add the negative pairs within iteration of unique hypotheses times unique premises.
@@ 5.1 Research Questions
5.2 Faithfulness Metrics

### v2

@@ Abstract
Ensuring complex systems meet regulations typically requires checking the validity of assurance cases through a claim-argument-evidence framework. Some challenges in this process include the complicated nature of legal and technical texts, the need for the model explanations, and limited access to assurance case data. We propose a compliance detection approach based on Natural Language Inference (NLI): EXplainable CompLiance detection with Argumentative Inference of Multi-hop reasoning (EXCLAIM). We formulate the claim-argument-evidence structure of an assurance case as a multi-hop inference for explainable and traceable compliance detection. We address the limited number of assurance cases by generating them using LLMs. We introduce metrics that measure the generated assurance cases’ coverage and structural consistency. We demonstrate the effectiveness of the generated assurance case from GDPR requirements in a multi-hop inference task as a case study. Our results highlight the potential of NLI-based approaches in automating the regulatory compliance process.
@@ 1 Introduction
Previous study (Cejas et al., 2023) uses inherently interpretable rule-based graph matching, but it still struggles with linguistic variability, ambiguous legal phrasing, and scalability to new regulations. A recent study (Azeem and Abualhaija, 2024) explores statistical machine learning and language model methods for the compliance detection task as text classification. However, it does not consider transparency and explainability factors such as model interpretability and reasoning.
To address the transparency and robustness factors of detecting risk and safety, experts use assurance cases (Rhodes et al., 2010; Mansourov and Campara, 2011) as a structured argument approach to audit complex software products in aircraft or autonomous vehicles to standard requirements. This is to ensure compliance with legal and technical standards. An assurance case is a systematic collection of arguments and supporting evidence to demonstrate a certain level of confidence that a product meets particular claims related to its requirements (Netkachova et al., 2015; Piovesan and Griffor, 2017). These assurance cases can be created and updated throughout the development lifecycle (Ross et al., 2022).
@@
One possible approach is to use deductive reasoning (Rushby et al., 2015) on the assurance case. Deductive reasoning in the assurance case derives conclusions from structured claims and arguments from evidence (Bloomfield and Bishop, 2009). This reasoning can be framed in the Natural Language Inference (NLI) task (Dagan et al., 2005; Bowman et al., 2015) to determine if a hypothesis follows from a premise. If there is a set of premises, multi-hop inference (Jansen, 2018; Yang et al., 2018) is needed for reasoning over interconnected claims and evidence. This makes multi-hop inference suitable for assurance cases that link high-level system properties to low-level evidence, like test results.
This paper proposes EXplainable CompLiance detection with Argumentative Inference of Multi-hop reasoning (EXCLAIM) based on Assurance Case. Figure 2 illustrates our workflow that explains this paper’s contribution, which provides a summary of this paper. From the figure, we utilise pre-trained transformer models (Devlin et al., 2019) to learn premise-hypothesis pairs for deductive multi-hop reasoning, enabling compliance detection. This reasoning process can detect non-compliance by providing a trace as an explanation of gaps between the main claim of fulfilled requirements and the evidence.
Specifically, our contributions address challenges in (i) explainable compliance detection, (ii) advanced reasoning and (iii) the lack of assurance case data as follows:
  • In terms of the model explainability, we analyse the faithfulness of interpretation methods applied to the NLI models for the compliance detection task. We consider four interpretation methods: gradient, Integrated gradients, LIME, and SHAP.
  • In terms of the data explainability and reasoning, we formulate the claim-argument-evidence structure from the assurance case as a discourse framework. This framework can be utilised to develop a multi-hop inference model for the explainable tracing of the compliance detection model’s output (Figure 1).
  • To address the lack of assurance case data, we propose evaluation methods to measure the assurance case generated by large language models (LLMs) to validate the reliability of instance-level consistency and structural coherence. We conduct a study by generating Claim Argument Evidence Assurance Case data with open source and proprietary LLMs from GDPR requirements as the context prompt from a previous study (Cejas et al., 2023). We analyse the generated Assurance Case data by using our proposed metrics.
@@ 4.1 GDPR-DPA
We use the compliance detection dataset from a previous study (Azeem and Abualhaija, 2024) for the NLI task on a real-world human-annotated dataset [^1] https://zenodo.org/records/11047441. We called this dataset the GDPR-DPA from the General Data Protection Regulation-Data Processing Agreement (GDPR-DPA). The details of the dataset statistics are presented in Table 2. The number of instances in the training set is larger than the original data due to pairing of negative samples between non-matching premises (requirements) and the hypothesis (DPA). We set the sampling rate to 0.1 to add negative pairs within the iteration of unique hypotheses times the number of unique premises.
@@ 5.1 Research Questions
5.2 Evaluating Explanation – Faithfulness Metrics
@@ 10 Ethical Considerations
11 Acknowledgements
This work has been funded by the European Commission under grant agreement No. 101120606, CERTIFAI. We thank Iker Lasa Ojanguren and Jose Arias Marin from FUNDACION TECNALIA RESEARCH & INNOVATION for their valuable contributions to the discussion and technical knowledge for Generating Assurance Case with Large Language Models as part of the CERTIFAI project. We thank Sunanda Bose for the technical discussion on compliance detection. We would also like to thank Akriti Sharma for helping to improve the clarity of the paper’s writing. This work has also benefited from the Experimental Infrastructure for Exploration of Exascale Computing (eX3), which is financially supported by the Research Council of Norway under contract 270053.